According to health data, about 17% of people drink diet soda.  A random sample of 400 people was chosen and the number who drink diet soda determined.  Find the probability that   is less than 0.15

A.

0

B.

0.4576

C.

0.3251

D.

0.8563

E.

none of these
Answer #1

Solution given that p=0.17 n = 400 1-² = 1-0.17 = 0.83 Mp=0.17 6² - P (1-P) - 0.17X0.83 0.01878 D ie o.15) = ê , P-M 6 f 0.15